A committee of ten people must be selected froma group containing 8 males and 9 females. What is the probability that the committee will have more females than males.
Answer by edjones(8007)

9F+1M = (9C9)(8C1)=8
8F+2M = (9C8)(8C2)=252
7F+3M = (9C7)(8C3)=2016
6F+4M = (9C6)(8C4)=5880
-----------------------------Add
===================>8156
.
.
8156/17C10 = .4194
.
Ed